OpenStudy (anonymous):

B(3;6) . C (-4;3) = x-x1/x2-x1=y-y1/y2-y1 =????

OpenStudy (anonymous):

B(3;6) . C (-4;3) = X-X1/X2-X1=Y-Y1/Y2-Y1 =????

OpenStudy (mertsj):

\[\frac{x-3}{-4-3}=\frac{y-6}{3-6}\] \[\frac{x-3}{-7}=\frac{y-6}{-3}\] \[-3x+9=-7y+42\] \[3x-7y=-33\]

OpenStudy (anonymous):

y+6/3-6=???

OpenStudy (mertsj):

you typed y-y1/y2-y1 so that is what I typed. I substituted 6 for y1 and 3 for y2